OpenStudy (anonymous):

using the definition of a derivative find f(x) =-2/sqrt{x}

OpenStudy (anonymous):

\[f(x)= -2/ \sqrt{x}\]

OpenStudy (anonymous):

this will be a pain to write out but not hard to compute. you must use the definition yes?

OpenStudy (anonymous):

i know that part i am having problems with rationalizing the denominator

OpenStudy (anonymous):

you need to compute \[lim_{h->0}\frac{\frac{-2}{\sqrt{x+h}}+\frac{2}{\sqrt{x}}}{h}\]

OpenStudy (watchmath):

Good opportunity to practice your LaTeX satellite! :)

OpenStudy (anonymous):

go ahead watchmath.

OpenStudy (anonymous):

i am learning. eventually i will be fluent in it

OpenStudy (anonymous):

i know that part the step after that

OpenStudy (anonymous):

ok lets give me a break and at least factor out the -2, since this has nothing to do with the limit ok?

OpenStudy (anonymous):

just factor out by 2 thats all you had t say lol thanks so much

OpenStudy (anonymous):

i knew that i was just doubting myself

OpenStudy (anonymous):

i knew that i was just doubting myself

OpenStudy (anonymous):

so we will just write \[lim_{h->0}\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

all divided by h of course. we worry about that last.

OpenStudy (anonymous):

so now we rationalize

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}=\frac{\sqrt{x}-\sqrt{x+h}}{\sqrt{x}\sqrt{x+h}}\]

OpenStudy (anonymous):

yes multiply numerator and denominator by the conjugate of th enumerator.

OpenStudy (anonymous):

which of course is \[\sqrt{x}+\sqrt{x+h}\]

OpenStudy (anonymous):

when you do that the numerator will just be \[x-x+h=h\]

OpenStudy (anonymous):

and the denominator will be the product \[\sqrt{x} \sqrt{x+h}) (\sqrt{x}+\sqrt{x+h})\]

OpenStudy (anonymous):

sorry am not repsonding i was working it out thanks i got it

OpenStudy (anonymous):

thats ok i was typing. did you get the numerator correctly? it is just h

OpenStudy (anonymous):

yes the x's cancel

OpenStudy (anonymous):

the denominator is that ugly thing i wrote last. so you have in total \[\frac{h}{h(\sqrt{x}\sqrt{x+h})(\sqrt{x}+\sqrt{x+h}}\]

OpenStudy (anonymous):

yes they 'cancel' meaning they add to zero. and of course dividing by h means the h goes in the denominator

OpenStudy (anonymous):

the h cancels

OpenStudy (anonymous):

so that last ugly thing i wrote is what you get when you rationalize the numerator. yes now the h cancels.

OpenStudy (anonymous):

leaving \[\frac{1}{(\sqrt{x}\sqrt{x+h})(\sqrt{x}+\sqrt{x+h})}\]

OpenStudy (anonymous):

now you can replace h by 0 since you will not be dividing by 0.

OpenStudy (anonymous):

to get... \[\frac{1}{(\sqrt{x}\sqrt{x})(\sqrt{x}+\sqrt{x})}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[=\frac{1}{2x\sqrt{x}}\]

OpenStudy (anonymous):

oh damn i made a mistake early on the numerator was \[x-(x+h)=-h\] not \[x-x+h\]

OpenStudy (anonymous):

a very bush league mistake but easily rectified. just replace the 1 in the numerator by -1

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

so the correct answer is \[\frac{-1}{2x\sqrt{x}}\]

OpenStudy (anonymous):

then don't forget to multiply by the -2 at the end because we factored it out.

OpenStudy (anonymous):

so the "final answer" as we say is \[\frac{1}{x\sqrt{x}}\] sorry it took a while

OpenStudy (anonymous):

all steps clear?

OpenStudy (anonymous):

sorry my computer is freezing idk if its thes site or what

OpenStudy (anonymous):

sometime the site is funky. earlier today it certainly was.

OpenStudy (anonymous):

if you have a question about any step post and i will respond

OpenStudy (anonymous):

i understand thank you sooo much!